# FANformer: Improving Large Language Models Through Effective Periodicity Modeling

### v1

@@ C.2 Evaluation Setup
Appendix E Extended results of Section 4.4.4
@@
Appendix F Derivation of $\text{ATF}(\mathbf{X})=\text{Attention}(\text{FANLayer}^{\prime}(\mathbf{X}))$
[…]
$=\text{Attention}(\mathbf{X}_{F})\quad\text{{(matches standard attention: $\text{Attention}(\mathbf{Z})=\text{softmax}\left(\frac{\mathbf{Z}\mathbf{W}_{Q}(\mathbf{Z}\mathbf{W}_{K})^{\top}}{\sqrt{d_{k}}}\right)\mathbf{Z}\mathbf{W}_{V}$)}}$
[…]
Appendix G Comprehensive Experimental Details
G.1 Detailed training settings of FANformer
We train FANformer-1B using the ZeRO optimizer strategy (Rajbhandari et al., 2020) via PyTorch’s DDP framework (Li, 2018). Following OLMo Groeneveld et al. (2024), we use a constant global batch size of approximately 4M tokens (2048 instances, each with a sequence length of 2048 tokens). To improve throughput, we employ PyTorch’s amp module with the bfloat16 format. We employ the AdamW optimizer Loshchilov et al. (2019) for the model’s training process. The learning rate for all LLMs is set to 4.0e-4. We warm up the learning rate over 2000 steps ( 8B tokens) and then decay it in a cosine manner from there down to a tenth of the peak learning rate over the remainder of training. We employ FlashAttention Dao et al. (2022) to accelerate the model training and inference processes, leveraging its ability to optimize memory usage and computational efficiency. The total GPU computational cost for pre-training FANformer-1B amounts to approximately 4,700 GPU hours.
G.2 Detailed Setup for Section 4.2
@@ G.1 Detailed training settings of FANformer
Table 6: Model size and setup used for FANformer in Section 4.2, where Transformers follows the setups of previous work OLMo Groeneveld et al. (2024).
@@
G.3 Validation Set And Downstream Tasks
Following (Zhu et al., 2024), we use V2 Validation Sets, V3 Validation Sets, and Downstream tasks to evaluate our approach. The specific tasks included in V2 validation sets, V3 validation sets, and downstream tasks are listed in Table 7.
@@ G.3 Validation Set And Downstream Tasks
piqa (Bisk et al., 2020)
hellaswag (Zellers et al., 2019)
winogrande (Sakaguchi et al., 2020)
openbook_qa (Mihaylov et al., 2018)
sciq (Welbl et al., 2017)
arc_easy (Clark et al., 2018)
copa (Roemmele et al., 2011)
commitment_bank (De Marneffe et al., 2019)
mrpc (Dolan et al., 2005)
rte (Dagan et al., 2005)
sst2 (Socher et al., 2013)
G.4 Detailed Setup of Case-based and Rule-based Reasoning.
Following the work (Hu et al., 2024), we focus on binary operations that take two numbers, $a$ and $b$, as inputs. Denoting $c$ as the target label, the constructed datasets are in the form of $\mathcal{D}=\{((a_{i},b_{i}),c_{i})\}$ for two mathematical tasks: modular addition and linear regression. The two tasks are defined as follows:
@@
The work (Hu et al., 2024) employs the Leave-Square-Out method to evaluate the generalization ability of the Transformer. In this approach, a square test set is created to isolate the test samples from the training samples. For instance, consider the center of the square at $(a_{k},b_{k})$ with a side length of $l_{k}$. The square test set is defined as $\mathcal{T}_{k}=\{((a_{i},b_{i}),c_{i})\mid a_{k}-\frac{l_{k}}{2}\leq a_{i}\leq a_{k}+\frac{l_{k}}{2},b_{k}-\frac{l_{k}}{2}\leq b_{i}\leq b_{k}+\frac{l_{k}}{2}\}$, and all remaining samples from the training set. This division creates a "hole" in the center of the training set, which is more challenging for the model compared to a random split. Since there are no similar cases in the training set to aid the model in solving the problem, this method tests whether the model has truly learned the underlying rules. In the experiments of the work  (Hu et al., 2024), they found that Transformer-based models fail to generate correct answers for the test set in the "hole". Therefore, we use this method to assess the generalization ability of FANformer.
@@
Following the work (Hu et al., 2024), we apply the Leave-Square-Out method to each dataset. Specifically, we extract a square comprising 441 samples (from a total of approximately 10,000 samples) with a side length of 20 to form our test set, leaving the remainder as the training set. It is important to note that, despite removing a small portion of training samples, we ensure that all tokens present in the dataset appear in the training set. This precaution is to prevent the models from failing simply due to encountering unseen tokens. We then proceed to finetune Transformer and FANformer models using this specific training-test split for each dataset.

### v2

@@ Appendix E Computational Cost Analysis
Appendix G Extended results of Section 4.4.3
@@
Appendix H Case for Section 4.4.4
We present a case of FANformer and the baselines under a logical reasoning stress-test in Figure 13.
Figure 13: Case of FANformer and the baselines under logical reasoning stress-test.
Appendix I Formal Proof of $\text{ATF}(\mathbf{X})=\text{Attention}(\text{FANLayer}^{\prime}(\mathbf{X}))$
Proposition 1.
For an input sequence representation $\mathbf{X}\in\mathbb{R}^{n\times d}$, the ATF operator satisfies
$\mathrm{ATF}(\mathbf{X})\;=\;\mathrm{Attention}\!\bigl{(}\mathrm{FANLayer}^{\prime}(\mathbf{X})\bigr{)}.$
Proof.
By the definition of ATF via Eq. (3), we have
$\text{ATF}(\mathbf{X})=\text{softmax}\left(\frac{\mathbf{Q}_{F}\mathbf{K}_{F}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V}_{F}.$
Substituting $\mathbf{Q}_{F},\mathbf{K}_{F},\mathbf{V}_{F}$ from Eq. (2)) into Eq. (3) yields
$\text{ATF}(\mathbf{X})=\text{softmax}\left(\frac{(\mathbf{X}_{F}\mathbf{W}_{Q})(\mathbf{X}_{F}\mathbf{W}_{K})^{\top}}{\sqrt{d_{k}}}\right)(\mathbf{X}_{F}\mathbf{W}_{V})$
Matching to standard attention, for any input $\mathbf{Z}$, multi-head attention (single head shown for clarity) is defined as
$\text{Attention}(\mathbf{Z})=\text{softmax}\left(\frac{\mathbf{Z}\mathbf{W}_{Q}(\mathbf{Z}\mathbf{W}_{K})^{\top}}{\sqrt{d_{k}}}\right)(\mathbf{Z}\mathbf{W}_{V})$
Thus, we have
$\mathrm{ATF}(\mathbf{X})=\mathrm{Attention}(\mathbf{X}_{F}).$ (12)
Finally, substituting $\mathbf{X}_{F}=\mathrm{FANLayer}^{\prime}(\mathbf{X})$ via Eq. (1) into the above formula establishes the claim:
$\mathrm{ATF}(\mathbf{X})=\mathrm{Attention}\!\bigl{(}\mathrm{FANLayer}^{\prime}(\mathbf{X})\bigr{)}.$
Appendix J Comprehensive Experimental Details
J.1 Detailed training settings of FANformer
We train FANformer-1B using the ZeRO optimizer strategy [Rajbhandari et al., 2020] via PyTorch’s DDP framework [Li, 2018]. Following OLMo [Groeneveld et al., 2024], we use a constant global batch size of approximately 4M tokens (2048 instances, each with a sequence length of 2048 tokens). To improve throughput, we employ PyTorch’s amp module with the bfloat16 format. We employ the AdamW optimizer [Loshchilov and Hutter, 2019] for the model’s training process. The learning rate for all LLMs is set to 4.0e-4. We warm up the learning rate over 2000 steps ( 8B tokens) and then decay it in a cosine manner from there down to a tenth of the peak learning rate over the remainder of training. We employ FlashAttention [Dao et al., 2022] to accelerate the model training and inference processes, leveraging its ability to optimize memory usage and computational efficiency. The total GPU computational cost for pre-training FANformer-1B amounts to approximately 47,600 GPU hours.
J.2 Detailed Setup for Section 4.2
@@ Appendix I Formal Proof of $\text{ATF}(\mathbf{X})=\text{Attention}(\text{FANLayer}^{\prime}(\mathbf{X}))$
Table 8: Model size and setup used for FANformer in Section 4.2, where Transformers follows the setups of previous work OLMo [Groeneveld et al., 2024].
@@
J.3 Validation Set And Downstream Tasks
Following [Zhu et al., 2024], we use V2 Validation Sets, V3 Validation Sets, and Downstream tasks to evaluate our approach. The specific tasks included in V2 validation sets, V3 validation sets, and downstream tasks are listed in Table 9.
@@
piqa [Bisk et al., 2020]
hellaswag [Zellers et al., 2019]
winogrande [Sakaguchi et al., 2020]
openbook_qa [Mihaylov et al., 2018]
sciq [Welbl et al., 2017]
arc_easy [Clark et al., 2018]
copa [Roemmele et al., 2011]
commitment_bank [De Marneffe et al., 2019]
mrpc [Dolan and Brockett, 2005]
rte [Dagan et al., 2005]
sst2 [Socher et al., 2013]
J.4 Detailed Setup of Case-based and Rule-based Reasoning.
Following the work [Hu et al., 2024], we focus on binary operations that take two numbers, $a$ and $b$, as inputs. Denoting $c$ as the target label, the constructed datasets are in the form of $\mathcal{D}=\{((a_{i},b_{i}),c_{i})\}$ for two mathematical tasks: modular addition and linear regression. The two tasks are defined as follows:
@@
The work [Hu et al., 2024] employs the Leave-Square-Out method to evaluate the generalization ability of the Transformer [Dong et al., 2024c]. In this approach, a square test set is created to isolate the test samples from the training samples. For instance, consider the center of the square at $(a_{k},b_{k})$ with a side length of $l_{k}$. The square test set is defined as $\mathcal{T}_{k}=\{((a_{i},b_{i}),c_{i})\mid a_{k}-\frac{l_{k}}{2}\leq a_{i}\leq a_{k}+\frac{l_{k}}{2},b_{k}-\frac{l_{k}}{2}\leq b_{i}\leq b_{k}+\frac{l_{k}}{2}\}$, and all remaining samples from the training set. This division creates a "hole" in the center of the training set, which is more challenging for the model compared to a random split. Since there are no similar cases in the training set to aid the model in solving the problem, this method tests whether the model has truly learned the underlying rules. In the experiments of the work  [Hu et al., 2024], they found that Transformer-based models fail to generate correct answers for the test set in the "hole". Therefore, we use this method to assess the generalization ability of FANformer.
@@
Following the work [Hu et al., 2024], we apply the Leave-Square-Out method to each dataset. Specifically, we extract a square comprising 441 samples (from a total of approximately 10,000 samples) with a side length of 20 to form our test set, leaving the remainder as the training set. It is important to note that, despite removing a small portion of training samples, we ensure that all tokens present in the dataset appear in the training set. This precaution is to prevent the models from failing simply due to encountering unseen tokens. We then proceed to finetune Transformer and FANformer models using this specific training-test split for each dataset.
J.5 Assessing LLMs’ Proficiency in Capturing Inferential Rules
Analysis Setup
Following the experimental setup proposed by Wang et al. [2024], we adopt the ULogic framework to systematically evaluate LLMs on their ability to capture underlying inferential logic. Specifically, we leverage a curated probing subset comprising 1,104 diverse rules drawn from their rule base. These rules—manually verified by the original authors—span a range of lengths, polarities, and structural patterns, ensuring broad coverage and high quality. The evaluation is framed as a binary entailment classification task, where the model must determine whether a given rule expresses a valid logical entailment. We employ a two-shot Chain-of-Thought (CoT) prompting method [Wei et al., 2022], in which each input includes one correct and one incorrect example to minimize label bias. The model is prompted not only to make a binary judgment but also to justify its reasoning, with an appended instruction such as "and also explain why."
To further enhance the reliability of the evaluation, we incorporate the Law of Non-Contradiction [Priest et al., 2006], which posits that statements of the form “If X, then Y” and “If X, then not Y” cannot simultaneously be true. Accordingly, for each original rule, we construct a flipped version by negating its conclusion. A rule is considered correctly classified only if the model affirms the original rule as true and rejects the flipped version as false, as illustrated below. We evaluate the base model of FANformer-1B, OLMo-1B, and Qwen2.5-1.5B, as well as GPT-4 on the two most challenging levels of ULogic for stress-testing (i.e., Length 3 and Length 4).
If Premise, then Conclusion__\__original. True
If Premise, then Conclusion__\__flipped. False
Appendix K Limitations
Second, our work is orthogonal to the existing approaches for revising the attention mechanism, i.e., our work can seamlessly incorporate them, as verified in the derivation provided in Appendix I. There are numerous variants of attention mechanisms, as discussed in the related work (Section 5), such as Flash Attention [Dao et al., 2022], MQA [Shazeer, 2019], and MLA [DeepSeek-AI et al., 2024b]. In this work, we only incorporate Flash Attention for necessary acceleration, while leaving the exploration of other approaches for future work.